A tennis club charges an annual fee and a hourly fee for court time. One year, Tony played for 39 h and paid $384. Sandra played for 51 h and paid $456. Find the annual fee and the hourly fee.
Answer by nerdybill(7384)

A tennis club charges an annual fee and a hourly fee for court time. One year, Tony played for 39 h and paid $384. Sandra played for 51 h and paid $456. Find the annual fee and the hourly fee.
.
Let x = annual fee
and y = hourly fee
.
from:"One year, Tony played for 39 h and paid $384."
x + 39y = 384 (equation 1)
.
from:"Sandra played for 51 h and paid $456"
x + 51y = 456 (equation 2)
.
Our system of equations:
x + 51y = 456
x + 39y = 384
.
multiply bottom eq by -1 and add together:
x + 51y = 456
-x - 39y = -384
------------------
12y = 72
y = $6 per hour (hourly fee)
.
Plug above into equation 1 and solve for x:
x + 39y = 384
x + 39(6) = 384
x + 234 = 384
x = $150 per year (annual fee)